OpenStudy (anonymous):

PLEASE HELP! Format: y-y1=m(x-x1) (2,-1) m=-3

OpenStudy (goten77):

(2,-1) (x1,y1) y-(-1)=-3(x-2) just plugging it in now get it in a y= form

OpenStudy (anonymous):

y-(-1)=-3(x-2) what nest?

OpenStudy (anonymous):

????